# Emergence and Localisation of Semantic Role Circuits in LLMs

### v1

@@ Appendix F Full Results
We add the full results for Pythia-14m, Pythia-410M and LLaMA-1B.
@@ F.1 RQ1: Full Localisation Results
⟦t⟧ Pythia-14M Pythia-410M LLaMA-1B Role T-5 T-10 T-20 T-5 T-10 T-20 T-5 T-10 T-20 Beneficiary 0.465 0.769 0.988 0.345 0.58 0.906 0.4 0.657 0.939 Instrument 0.445 0.596 0.833 0.445 0.596 0.833 0.484 0.697 0.895 Location 0.368 0.557 0.841 0.368 0.557 0.841 0.413 0.636 0.882 Time 0.397 0.628 0.974 0.397 0.628 0.974 0.443 0.677 0.961
Table 8: Top-$K$ mass concentration at final checkpoint (143k steps). Values show fraction of total attribution mass captured by $K$ highest-mass nodes.
⟦t⟧ Role Pythia-14M Pythia-410M LLaMA-1B Beneficiary 11 / 13 / 15 16 / 20 / 23 14 / 17 / 22 Instrument 11 / 14 / 17 19 / 24 / 28 15 / 21 / 25 Location 12 / 15 / 17 19 / 23 / 27 17 / 21 / 25 Time 10 / 13 / 17 15 / 18 / 20 14 / 17 / 20
Table 9: Minimal node count $k$ required to capture 80% / 90% / 95% of total attribution mass at final checkpoint (143k steps).
@@
width=0.8⟧figures/cross_scale_spectral_beneficiary_pythia.png Refer to captionwidth=0.8⟧figures/cross_scale_spectral_location_pythia.png Refer to captionwidth=0.8⟧figures/cross_scale_spectral_instrument_pythia.png Refer to captionwidth=0.8⟧figures/cross_scale_spectral_time_pythia.png
Figure 6: Cross scale spectral distance
width=0.8⟧figures/cross_scale_overlap_beneficiary_pythia.png Refer to captionwidth=0.8⟧figures/cross_scale_overlap_location_pythia.png Refer to captionwidth=0.8⟧figures/cross_scale_overlap_instrument_pythia.png Refer to captionwidth=0.8⟧figures/cross_scale_overlap_time_pythia.png
Figure 7: Cross scale spectral distance
@@ F.2 RQ2: Full Emergence Dynamics
Two-segment piecewise fits to Top-$K$ mass trajectories yield very wide bootstrap confidence intervals (e.g. Time: ⟦512, 10k⟧; Instrument: ⟦64, 8k⟧), confirming that structural sparsification is gradual rather than concentrated at a discrete transition. Visual inspection of sparsity curves reveals smooth monotonic growth without identifiable inflection points. The structural substrate evolves continuously even when functional utility displays sharp changes.
@@
By convergence (143k steps), all circuits are highly compact: Top-20 mass ranges from 0.83–0.97, and only 20–28 nodes suffice to cover 95% of attribution. Time is the most concentrated (95% mass in 20 nodes), whereas Instrument and Location have slightly broader but still compact top-tiers.
@@
Across all roles in Pythia–14M, emergence is a gradual process in which structural properties stabilise early and monotonically, while functional utility develops in a noisy, role-dependent manner. Indispensability can lag far behind consolidation (e.g. 70k-step lag for Time), indicating that circuits may be structurally “pre-allocated” long before the model consistently relies on them. Together, these results support the conclusion that semantic-role circuits do not undergo discrete phase transitions but instead emerge through continuous refinement shaped by heterogeneous task signals and optimisation dynamics.
@@ F.3 Additional Semantic Roles
In addition to the four core roles analysed in the main paper, we applied the COMPASS pipeline to four further predicate–argument relations frequently used in semantic role labelling, including Path, Source, and Topic. Figure 9 reports their emergence trajectories (faithfulness, density, Top-$K$ mass) across training checkpoints.
Overall, these supplementary roles exhibit the same qualitative patterns identified for the main roles. Structural metrics (density and Top-$K$ mass) increase smoothly and monotonically, consistent with gradual sparsification rather than discrete phase transitions. In contrast, faithfulness trajectories are highly variable: Path shows an extremely sharp early increase, becoming highly faithful almost immediately, followed by stable mid-training behaviour. Source rises early but then gradually declines, mirroring the non-monotonic pattern seen in Time. Topic increases sharply at early steps before collapsing and then partially recovering, indicating heavy competition with other predictive cues. The extended Goal set displays declining faithfulness after an early peak, despite stable and increasing structural sparsity. These patterns match the dissociation seen in the primary roles: structural sparsity is stable and monotonic, but functional utility is noisy and task-pressure dependent. As before, structure often stabilises long before a role becomes functionally useful.
⟦h!⟧ Refer to captionwidth=⟧figures/emergence_curves.png
@@
⟦h!⟧ Refer to captionwidth=⟧figures/emergence_curves_1b_full.png
Figure 9: Emergence trajectories for supplementary semantic roles for the 1B model. Each row corresponds to a role; columns show faithfulness, density, and Top-$K$ mass across training.
⟦t⟧ Role $t_{\mathrm{ind}}$ (steps) $t_{\mathrm{cons}}$ (steps) det. threshold Goal 0 1,000 1.021 Path 8 0.803 Source 0 8 1.215 Topic 512 1,000 1.168
Table 10: Indispensability $(t_{\mathrm{ind}})$, and consolidation $(t_{\mathrm{cons}})$ per role for 1B model.
⟦t⟧ Role $\hat{t}_{c}$ (steps) 95% CI Goal 128 ⟦128, 7,000⟧ Path 1,000 ⟦8, 10,000⟧ Source 1,000 ⟦8, 10,000⟧ Topic 1,000 ⟦8, 10,000⟧
Table 11: Change-point estimates for Top-$K$ node mass via two-segment piecewise linear regression.
@@ F.4.1 Architectural Stratification at Convergence
⟦h!⟧ Type Role Nodes Heads Architecture 1 Time 16 0 Pure MLP chain; completely attention-free Path 18 2 Predominantly MLP with early framing (a0.h6) and late value composition (a12.h2+V) 2 Goal 18 3 Mid-layer extraction (a6.h0) with late integration (a12.h2+V, a14.h5) Topic 19 4 Early frame detection (a0.h2), mid-layer (a6.h0), late integration (a12.h2, a14.h5+V) Source 19 5 Rich multi-stage: early (a3.h7), mid (a6.h0), late (a12.h2, a14.h5, a15.h5) 3 Instrument 20 4 Mid-to-late extraction and integration (a3.h7, a6.h0, a12.h2, a14.h5) Location 20 4 Distributed integration (a1.h0, a12.h2, a13.h4, a14.h5) 4 Beneficiary 22 6 Complex late-stage architecture (a1.h0, a6.h0, a12.h2, a14.h5+V, a15.h0, a15.h5)
@@
Time and Path rely predominantly on MLP computation with minimal or zero attention. Time achieves the most extreme compression (16 nodes, zero attention), implementing a pure feed-forward pathway. Pathexhibits minimal attention involvement (18 nodes, 2 heads), we hypothesise a0.h6 provides early syntactic composition, while a12.h2 performs targeted value composition for the final pre-logit refinement. The dominant causal flow remains through the MLP backbone, with attention providing what seems to be an auxiliary support. Both roles mark a highly regular, closed set of expressions (e.g., temporal markers, path markers), enabling direct lexical classification without extensive need for contextual integration.
@@
Beneficiary has the most elaborate architecture at convergence (22 nodes, 6 heads), maintaining rich late-stage connectivity with dual final-layer heads. The circuit recruits a1.h0 for early processing, a6.h0 for mid-layer extraction, and four late-stage heads: a12.h2 for primary integration, a14.h5+V for feature extraction, and two heads (a15.h0, a15.h5) for final disambiguation. This architecture suggests Beneficiary requires parallel processing pathways to resolve persistent ambiguities, and probably the model maintains alternative hypotheses until the last computation step. The circuit’s relative complexity (22 nodes vs. 16–20 for other roles) and late-stage density indicate that beneficiary marking, despite being syntactically constrained (e.g., “for $\dots$”), requires more elaborate compositional reasoning than other participant roles, likely to distinguish benefactive readings from alternative interpretations.
@@ F.4.2 Developmental Trajectories Across Training
Circuit evolution from initialisation (step 32) to convergence (step 143k) reveals role-specific developmental patterns. Figures 10, 11 and 12 present a sample of causal-flow diagrams at three key checkpoints, early training (step 32), mid-training (step 71000), and convergence (step 143000). These snapshots expose systematic differences in how semantic-role circuits emerge and stabilise. We study them for all roles below.
⟦ht⟧ Refer to captionwidth=0.6⟧figures/time_causal_flow.png
@@
⟦ht⟧ Refer to captionwidth=0.6⟧figures/location_causal_flow.png
@@
⟦ht⟧ Refer to captionwidth=0.6⟧figures/Instrument_causal_flow.png
@@
Roles follow distinct consolidation trajectories aligned with semantic complexity. What we assume are more Template-matching roles (Time, Path) eliminate most attention: Time achieves pure MLP architecture by mid-training; Path retains minimal heads (a0.h6, a12.h2+V). Integration-dependent roles (Goal, Instrument, Location) stabilise core architectures by mid-training with only minor late refinement, converging to 2–4 attention heads with mid-to-late integration. Multi-phase processing roles (Topic, Source, Beneficiary) undergo non-monotonic reorganisation: Topic contracts then reinstates early framing (a0.h2) and late value composition (a14.h5+V); Source adds a3.h7 only at convergence; Beneficiary maintains persistent complexity with six heads including unique dual layer-15 architecture (a15.h0, a15.h5) for parallel disambiguation.
@@ F.5 Paraphrase-Based Within-Role Scaffold Controls (Pythia–1B)
⟦h!⟧ Role Faithfulness Top–20 Mass Instrument 0.657(+0.021)0.657\;(+0.021) 0.925(+0.009)0.925\;(+0.009) Location 0.731(0.053)0.731\;(-0.053) 0.920(+0.023)0.920\;(+0.023)
@@
⟦ht⟧ Method Faithful Path-spec. Granularity Scalable Notes / Risks Linear probing $\circ$ $\times$ token/residual $\checkmark$ Correlational, not causal Attribution Patching (EAP) (nanda2023attribution; syed-etal-2024-attribution) $\circ$ $\checkmark$ head/MLP edge $\checkmark$ Gradient-based; false negatives AtP (kramar2024atp) $\checkmark$ $\checkmark$ head/MLP edge $\checkmark\checkmark$ improved faithfulness; residual false negatives Temporal EAP-IG* $\checkmark$ $\checkmark$ head/MLP edge $\checkmark\checkmark$ Baseline/path sensitivity Path patching (GoldowskyDill2023LocalizingMB) $\checkmark$ $\checkmark\checkmark$ head/MLP path $\circ$ Expensive over checkpoints Causal scrubbing (chan2022causal) $\checkmark\checkmark$ $\checkmark\checkmark$ hypothesis-level $\times$ High implementation cost SAE (templeton2024scaling) $\checkmark$ $\checkmark$ feature-level $\circ$ SAE training; feature drift Transcoders (dunefsky2024transcoders) $\checkmark$ $\checkmark$ MLP sublayer $\circ$ Surrogate training; not causal Circuit tracing / attribution graphs (ameisen2025circuit) $\circ$ $\checkmark\checkmark$ feature–feature $\times$ Surrogate fidelity; prompt-specific

### v2

@@ Appendix F Full Results
We add the full results for Pythia-14m, Pythia-410M and LLaMA-1B. We note that all experiments were repeated on five different random seeds, and the reported results are the averaged graphs per model.
@@ F.1 RQ1: Full Localisation Results
Pythia-14M Pythia-410M LLaMA-1B
Role T-5 T-10 T-20 T-5 T-10 T-20 T-5 T-10 T-20
Beneficiary 0.465 0.769 0.988 0.345 0.58 0.906 0.4 0.657 0.939
Instrument 0.445 0.596 0.833 0.445 0.596 0.833 0.484 0.697 0.895
Location 0.368 0.557 0.841 0.368 0.557 0.841 0.413 0.636 0.882
Time 0.501 0.713 0.985 0.445 0.610 0.852 0.443 0.677 0.961
Table 8: Top-$K$ mass concentration at final checkpoint (143K steps). Values show fraction of total attribution mass captured by $K$ highest-mass nodes, demonstrating strong localisation across roles and model scales.
Role Pythia-14M Pythia-410M LLaMA-1B
Beneficiary 11 / 13 / 15 16 / 20 / 23 14 / 17 / 22
Instrument 11 / 14 / 17 19 / 24 / 28 15 / 21 / 25
Location 12 / 15 / 17 19 / 23 / 27 17 / 21 / 25
Time 10 / 14 / 18 18 / 24 / 28 14 / 17 / 20
Table 9: Minimal node count required for coverage at the final checkpoint (143K steps). Entries report the smallest number of nodes ($k$) needed to capture 80%, 90%, and 95% of total attribution mass for each role and model, where small number of nodes constitute the majority of mass.
@@
Figure 6: Cross-scale spectral distance of role circuits. Heatmaps show pairwise spectral distances between role circuits across Pythia model scales (14M, 410M, 1B), computed from the Top-50 edges and the first 20 eigenvalues. Lower values between larger models indicate greater similarity in information-flow geometry despite possible topological differences.
Figure 7: Cross-scale overlap of high-importance components. Heatmaps report average Top-30 node overlap (left) and Top-30 edge overlap (right) between role circuits across Pythia model scales. Node overlap consistently exceeds edge overlap, indicating reuse of key components with divergent routing across scales.
@@ F.2 RQ2: Full Emergence Dynamics
Two-segment piecewise fits to Top-$K$ mass trajectories yield very wide bootstrap confidence intervals (e.g. Time: [54, 5k]; Instrument: [64, 8K]), confirming that structural sparsification is gradual rather than concentrated at a discrete transition. Visual inspection of sparsity curves reveals smooth monotonic growth without identifiable inflection points. The structural substrate evolves continuously even when functional utility displays sharp changes.
@@
By convergence (143K steps), all circuits are highly compact: Top-20 mass ranges from 0.83–0.98, and only 15–18 nodes suffice to cover 95% of attribution. Beneficiary is the most concentrated (95% mass in 15 nodes), whereas Instrument and Location have slightly broader but still compact top-tiers.
@@
Across all roles in Pythia–14M, emergence is a gradual process in which structural properties stabilise early and monotonically, while functional utility develops in a noisy, role-dependent manner. Indispensability can lag far behind consolidation, indicating that circuits may be structurally “pre-allocated” long before the model consistently relies on them. Together, these results support the conclusion that semantic-role circuits do not undergo discrete phase transitions but instead emerge through continuous refinement shaped by heterogeneous task signals and optimisation dynamics.
@@ F.3 Additional Semantic Roles
In addition to the four core roles analysed in the main paper, we applied the COMPASS pipeline to four further predicate–argument relations frequently used in semantic role labelling, including Path, Source, and Topic. Figure 9 reports their emergence trajectories (faithfulness, density, Top-$K$ mass) across training checkpoints, and Tables 10, 11, and 12 demonstrate the indispensability, concentrations and change-point estimation for them, respectively.
Overall, these supplementary roles exhibit comparable qualitative patterns identified for the main roles. Structural metrics (density and Top-$K$ mass) increase smoothly and monotonically, consistent with gradual sparsification rather than discrete phase transitions. In contrast, faithfulness trajectories are highly variable: Path exhibits high volatility: faithfulness increases through early training, then spikes mid-training (step 71k) before collapsing back, indicating heavy competition with other predictive cues. Source shows high faithfulness early training, the declines and remains consistently throughout training despite structural consolidation. Topic displays an early spike before settling into moderate stability. Goal maintains moderate faithfulness with fluctuations mid-training, never achieving the stability seen in others. These patterns match the dissociation seen in the primary roles: structural sparsity is stable and monotonic, but functional utility is noisy and task-pressure dependent. As before, structure often stabilises long before a role becomes functionally useful.
@@
Figure 9: Emergence trajectories for supplementary semantic roles for the Pythia-1B model. Each row corresponds to a role; columns show faithfulness, density, and Top-$K$ mass across training.
Role $t_{\mathrm{ind}}$ (steps) $t_{\mathrm{cons}}$ (steps)
Goal 0 1,000
Path 64 1,000
Source 512 5,000
Topic 512 5,000
Table 10: Indispensability and consolidation timings for additional roles (Pythia-1B model). The table reports the earliest step at which each role becomes indispensable ($t_{\mathrm{ind}}$), and the step at which structural consolidation occurs ($t_{\mathrm{cons}}$)
Role Top-5 Top-10 Top-20 $k$ for 80/90/95%
Goal 0.507 0.708 0.931 14 / 18 / 22
Path 0.348 0.594 0.889 16 / 21 / 25
Source 0.421 0.655 0.917 15 / 19 / 23
Topic 0.444 0.657 0.908 15 / 20 / 25
Table 11: Final-step concentration of additional role circuits. Similar to results in the main paper, a small $k$ values show a limited subset of nodes constituent the majority part of computation.
Role $\hat{t}_{c}$ (steps) 95% CI
Goal 10,000 [128, 10,000]
Path 10,000 [128, 128,000]
Source 5,000 [128, 99,000]
Topic 5,000 [128, 99,000]
Table 12: Change-point estimates for Top-$K$ node mass. Estimates are obtained via two-segment piecewise linear regression, with 95% bootstrap confidence intervals shown for each role; wide intervals indicate gradual structural evolution.
@@ F.4.1 Architectural Stratification at Convergence
Type Role Nodes Heads Architecture
1 Path 18 2 Predominantly MLP with early framing (a0.h6) and late value composition (a12.h2+V)
Goal 18 3 Mid-layer extraction (a6.h0) with late integration (a12.h2+V, a14.h5)
2 Topic 19 4 Early frame detection (a0.h2), mid-layer (a6.h0), late integration (a12.h2, a14.h5+V)
Source 19 5 Rich multi-stage: early (a3.h7), mid (a6.h0), late (a12.h2, a14.h5, a15.h5)
3 Instrument 20 4 Mid-to-late extraction and integration (a3.h7, a6.h0, a12.h2, a14.h5)
Location 20 4 Distributed integration (a1.h0, a12.h2, a13.h4, a14.h5)
4 Beneficiary 22 6 Complex late-stage architecture (a1.h0, a6.h0, a12.h2, a14.h5+V, a15.h0, a15.h5)
Time 22 7 Diverse multi-stage architecture with integration of several components (a3.h7, a3.h5, a6.h0, a9.h2, a10.h1, a12.h2+V, a14.h5+V, a15)
@@
Goal and Path rely predominantly on MLP computation with minimal or zero attention. Path exhibits minimal attention involvement (18 nodes, 2 heads), we hypothesise a0.h6 provides early syntactic composition, while a12.h2 performs targeted value composition for the final pre-logit refinement. The dominant causal flow remains through the MLP backbone, with attention providing what seems to be an auxiliary support. Goal (18 nodes, 3 heads) implements a clean two-stage architecture: mid-layer feature extraction at a6.h0, followed by late-stage integration at a12.h2+V (with value composition edge) and final refinement at a14.h5. The circuit exhibits sparse connectivity with clear information bottlenecks at m6 and m12. Both roles mark a highly regular, closed set of expressions (e.g., path markers), enabling direct lexical classification without extensive need for contextual integration.
@@
Beneficiary and Time exhibit the most elaborate architectures at convergence (22 nodes each, 6–7 attention heads), maintaining rich late-stage connectivity with distributed processing across multiple layers. The circuit recruits a1.h0 for early processing, a6.h0 for mid-layer extraction, and four late-stage heads: a12.h2 for primary integration, a14.h5+V for feature extraction, and two heads (a15.h0, a15.h5) for final disambiguation. This architecture suggests Beneficiary requires parallel processing pathways to resolve persistent ambiguities, and probably the model maintains alternative hypotheses until the last computation step. The circuit’s relative complexity (22 nodes vs. 16–20 for other roles) and late-stage density indicate that beneficiary marking, despite being syntactically constrained (e.g., “for $\dots$”), requires more elaborate compositional reasoning than other participant roles, likely to distinguish benefactive readings from alternative interpretations. Time (22 nodes, 7 heads) exhibits similarly complex multi-stage processing with attention distributed across early (a3.h7), mid-layer (a6.h0, a9.h2, a10.h1), and late integration (a12.h2, a14.h5+V, a15). The circuit shows convergent information flow through mid-layer MLPs (m6, m9, m10) to late-stage integration nodes (m12, m14), with value composition at a14.h5 enabling feature extraction for temporal disambiguation. This architectural complexity is surprising given that temporal expressions often involve closed-class markers (“during the”, “at the”), but likely reflects the need to distinguish temporal from locative interpretations of ambiguous scaffolds (“at the”) and to resolve context-dependent temporal reference.
@@ F.4.2 Developmental Trajectories Across Training
Circuit evolution from initialisation to convergence (step 143K) reveals role-specific developmental patterns. Figures 10, 11 and 12 present a sample of causal-flow diagrams at three key checkpoints, early training (step 32), mid-training (step 71000), and convergence (step 143000). These snapshots expose systematic differences in how semantic-role circuits emerge and stabilise. We study them for all roles below.
@@
Roles follow distinct consolidation trajectories aligned with semantic complexity. What we assume are more Template-matching roles (Goal, Path) eliminate most attention: Path retains minimal heads (a0.h6, a12.h2+V). Integration-dependent roles ( Instrument, Location) stabilise core architectures by mid-training with only minor late refinement, converging to 2–4 attention heads with mid-to-late integration. Multi-phase processing roles (Topic, Source, Beneficiary) undergo non-monotonic reorganisation: Topic contracts then reinstates early framing (a0.h2) and late value composition (a14.h5+V); Source adds a3.h7 only at convergence; Beneficiary and Time maintain persistent complexity with many heads likely required for parallel disambiguation.
@@ F.5 Paraphrase-Based Within-Role Scaffold Controls (Pythia–1B)
Metric n Pearson $r$ p-value Spearman $\rho$
Circuit Size (Nodes) 8 -0.066 0.876 0.049
Attention Heads 8 -0.283 0.497 -0.268
Sparsity (Top-20 Mass) 8 0.296 0.476 0.311
Concentration (Gini) 8 0.326 0.430 0.357
Consolidation Time ($t_{cons}$) 8 -0.054 0.899 0.062
Table 14: Correlations between role frequency (sample size) and circuit properties at convergence (Pythia-1B, step 143K). All correlations show, indicating weak relationships between role frequency and circuit properties. Sample sizes range from 491 to 1,212 examples per role.
Role Faithfulness Top–20 Mass
Instrument 0.657(+0.021)0.657\;(+0.021) 0.925(+0.009)0.925\;(+0.009)
Location 0.731(0.053)0.731\;(-0.053) 0.920(+0.023)0.920\;(+0.023)
@@
Circuit Property n Pearson r p-value Spearman $\rho$
Circuit Size (nodes) 8 0.346 0.401 0.439
Attention Heads 8 0.310 0.455 0.244
Sparsity (Top-20 mass) 8 0.106 0.803 0.095
Concentration (Gini) 8 0.430 0.288 0.381
Table 16: Corpus frequencies from 5,000 documents sampled from The Pile and the correlations between corpus scaffold frequency and circuit properties at convergence (Pythia-1B, step 143K). All correlations show $|r|\leq 0.430$, $p\geq 0.288$ (not significant).
Method Faithful Path-spec. Granularity Scalable Notes / Risks
Linear probing $\circ$ $\times$ token/residual $\checkmark$ Correlational, not causal
Attribution Patching (EAP) (nanda2023attribution; syed-etal-2024-attribution) $\circ$ $\checkmark$ head/MLP edge $\checkmark$ Gradient-based; false negatives
AtP (kramar2024atp) $\checkmark$ $\checkmark$ head/MLP edge $\checkmark\checkmark$ improved faithfulness; residual false negatives
Temporal EAP-IG* $\checkmark$ $\checkmark$ head/MLP edge $\checkmark\checkmark$ Baseline/path sensitivity
Path patching (GoldowskyDill2023LocalizingMB) $\checkmark$ $\checkmark\checkmark$ head/MLP path $\circ$ Expensive over checkpoints
Causal scrubbing (chan2022causal) $\checkmark\checkmark$ $\checkmark\checkmark$ hypothesis-level $\times$ High implementation cost
SAE (templeton2024scaling) $\checkmark$ $\checkmark$ feature-level $\circ$ SAE training; feature drift
Transcoders (dunefsky2024transcoders) $\checkmark$ $\checkmark$ MLP sublayer $\circ$ Surrogate training; not causal
Circuit tracing / attribution graphs (ameisen2025circuit) $\circ$ $\checkmark\checkmark$ feature–feature $\times$ Surrogate fidelity; prompt-specific
@@
F.6 Role Frequency and Circuit Properties
F.6.1 Sample Size and Circuit Properties
To test whether circuit architecture is determined by the availability of samples in our filtered dataset, we examined correlations between sample sizes and circuit properties at convergence (Pythia-1B, step 143K). If circuit complexity were driven primarily by the volume of valid examples, we would expect positive correlations between sample sizes and measures of circuit size or structural complexity.
Correlation Analysis.
Table 14 reports Pearson and Spearman correlations between filtered sample sizes and five circuit properties: circuit size (nodes), attention head count, sparsity (Top-20 mass), concentration (Gini coefficient), and consolidation timing ($t_{cons}$). All correlations show $|r|<0.35$ with all $p>0.4$, indicating negligible relationships between sample size and circuit architecture. The strongest correlation (Gini: $r=0.326$, $p=0.430$) remains statistically
Role-Type Stratification.
In contrast to the absence of sample-size effects, circuit architecture stratifies clearly by role type (Table 13). Roles requiring complex disambiguation (Type 4: BENEFICIARY, TIME) develop the largest circuits (22 nodes, 6–7 attention heads) independent of their sample sizes (621 vs. 753 examples). Conversely, lexical pattern matching (Type 1: PATH) converges to a minimal architecture (18 nodes, 2 heads) despite moderate sample availability (707 examples). Multi-stage compositional roles (Type 2) and balanced hybrid architectures (Type 3) occupy intermediate positions, with circuit complexity determined by semantic processing demands rather than example count.
The consistent absence of correlations between sample sizes and all circuit properties demonstrates that architectural complexity is not determined by the number of valid examples in our filtered dataset.
F.6.2 Scaffold Frequency in Training Corpus
To validate that the selected semantic roles are well-represented in the training data and to test whether circuit architecture can be explained by training signal strength alone, we analysed scaffold frequencies in the Pythia training corpus (The Pile (pile)). We sampled 5000 documents ($\sim$50M tokens) using stratified sampling (every 4,200th document) to ensure representative coverage across Pile subsets, and counted occurrences of the prepositional scaffolds used in our role-conditioned continuation task. We then correlated these corpus frequencies with circuit properties at convergence to test whether training frequency determines architectural complexity.
Corpus Frequencies.
All roles are well-represented in the training data, ranging from 517.9 (Topic) to 5,498.9 (Location) instances per million tokens. Locative and temporal roles exhibit the highest frequencies (Location: 5,498.9/1M; Time: 5,452.3/1M), reflecting the prevalence of spatial and temporal modification in natural language. Directional roles show intermediate frequencies (Goal: 4,147.6/1M; Source: 1,057.2/1M; PATH: 598.2/1M), while participant and propositional roles appear less frequently but remain well represented (Instrument: 1,990.5/1M; Beneficiary: 1,523.3/1M; TOPIC: 517.9/1M).
Correlation with Circuit Architecture.
Table 16 reports Pearson and Spearman correlations between corpus scaffold frequencies and circuit properties at convergence (step 143K). All correlations are weak to moderate ($|r|\leq 0.430$) and statistically non-significant ($p>0.28$), indicating that training frequency does not reliably predict circuit architecture. The strongest correlation (Gini coefficient: $r=0.430$, $p=0.288$) explains only 18% of variance and is not statistically significant. Circuit size and attention head count show weak positive trends ($r=0.346$ and $r=0.310$ respectively), while sparsity (Top-20 mass) shows virtually no relationship ($r=0.106$, $p=0.803$). We note that one of these relationships approaches statistical significance at $\alpha=0.05$ level.
The absence of systematic frequency effects is most clearly demonstrated by roles with nearly identical corpus frequencies but divergent architectures. Location and Time appear at virtually the same frequency in the training data (5,498.9 vs. 5,452.3 per million tokens, a difference of 0.8%), but develop different circuit structures. Location converges to a balanced hybrid architecture (20 nodes, 4 attention heads) with distributed integration across layers (a1.h0, a12.h2, a13.h4, a14.h5), while Time develops a more complex multi-stage architecture (22 nodes, 7 attention heads) with rich attention involvement distributed across early (a3.h7), mid-layer (a6.h0, a9.h2, a10.h1), and late integration stages (a12.h2, a14.h5, a15). This architectural difference, despite a near-identical training signal, demonstrates that circuit organisation cannot be predicted from corpus frequency alone. The results indicate that instead of building circuits proportional to how often each role appears in training, models develop architectures tailored to the semantic processing demands of each role. This suggests that circuit organisation reflects functional requirements, the computational work needed to bind each semantic role, rather than statistical regularities in the training corpus.